Evaluate the discriminant of each equation. How many real and imaginary
solutions does each have?(please help me with this.. i realy need it done in like a very fast time.. please please))
37. x2 + 5x + 6 = 0
38. 3x2 - 4x + 3 = 0
39. -2x2 - 5x + 4 = 0
40. 16x2 - 8x + 1 = 0

Answer by edjones(8007)

ax^2+bx+c=0
The dicriminant is b^2-4ac.
If the discriminant is negative then the solutions will be imaginary. If it is zero then there will be only one real solution.
37. x2 + 5x + 6 = 0
a=1, b=5, c=6
5^2-(4*6) is the discriminant.
25-24=1 two real solutions.
.
38. 3x2 - 4x + 3 = 0
a=3 b=-4 c=3
(-4)^2-(4*3*3) is the discriminant
16-36=-20 two imaginary solutions
.
39. -2x2 - 5x + 4 = 0
a=-2 b=-5 c=4
(-5)^2-(4*-2*4) is the discriminant.
25-32=-7 two imaginary solutions
.
40. 16x2 - 8x + 1 = 0
a=16, b=-8, c=1
(-8)^2-(4*16*1)
64-64=0 one real solution
.
Ed